Operating and Cash Cycles.

ABC firm practices just-in-time inventory management and carries a negligible amount of inventory. Its accounts receivable turnover is 6 times and its accounts payable turnover is 4 times. Assuming that the year has 360 days, Compute the operating and cash cycles for ABC firm and discuss its implications on its working capital management.
Answer #1

Operating cash cycle = Days of Inventory + Days of accounts receivables

The company has just in time inventory management system. Hence there is no inventory in hand at any time and inventory period = 0

Accounts receivable days = 360 / 6 times = 60 days

Operating cash cycle = 0 + 60 = 60 days

Cash cycle = Operating cash cycle - Accounts payable period

Accounts payable period = 360 / accounts payable turnover = 360 / 4 = 90 days

Cash cycle = 60 -90 = - 30 days